OpenStudy (anonymous):

Definite Integral: integrate 1 to e (x^2+x+1)/(x) dx

OpenStudy (fibonaccichick666):

ewwy... hmmm let's see what we can do

OpenStudy (fibonaccichick666):

first, is this the question: \[\int_1^e \frac{x^2+x+1}{x} dx\]

OpenStudy (fibonaccichick666):

@ZeroAT

OpenStudy (freckles):

hint: (a+b)/c=a/c+b/c

OpenStudy (anonymous):

\[\int\limits_{1}^{e} \frac{ x^2+x+1 }{ x } \rightarrow x + 1 + \frac{ 1 }{ x } \rightarrow \frac{ x^2 }{ 2 } + x + \ln(x) + C\] F(e) – F(1) gives the final result of e²/2 + e – 1/2 Thanks for the hint freckles!

OpenStudy (fibonaccichick666):

yep that's very close. check your lower bounds result though

OpenStudy (freckles):

i think he combined like terms if i didn't make a mistake that seems to be right (also good job; some people would be looking at the hint all day )

OpenStudy (fibonaccichick666):

ah ln e=1, yes you are correct freckles